# CLIX: Cross-Lingual Explanations of Idiomatic Expressions

### v2

@@ Abstract
Automated definition generation systems have been proposed to support vocabulary expansion for language learners. The main barrier to the success of these systems is that learners often struggle to understand definitions due to the presence of potentially unfamiliar words and grammar, particularly when non-standard language is involved. To address these challenges, we propose CLIX, the task of Cross-Lingual explanations of Idiomatic eXpressions. We explore the capabilities of current NLP models for this task, and observe that while it remains challenging, large language models show promise. Finally, we perform a detailed error analysis to highlight the key challenges that need to be addressed before we can reliably incorporate these systems into educational tools.
@@ 1 Introduction
Motivated by these challenges, we propose the task of Cross-Lingual explanation of Idiomatic eXpressions (CLIX). Given a source language idiomatic expression, we aim to generate a natural language explanation in a specified target language. As opposed to definition, we use the term explanation to allow for more flexibility in tailoring explanations to certain situations as well as potential inclusions of context. We hypothesize that obtaining explanations of idiomatic expressions in the learner’s first language removes many of the barriers to understanding introduced by traditional definition generation systems. We choose to focus on idiomatic expressions as they are an important element of language learning that is particularly challenging for learners and automated systems alike. Consider the utterance, he and I don’t see eye to eye on a variety of topics. The idiomatic expression contained within this sentence is not composed of particularly challenging words, yet in this instance it denotes a negative stance relative to the referent’s opinions and its meaning has nothing to do with the general definitions of the words see and eye. This example helps to display the fact that idiomatic expressions often are of unpredictable form, appear in limited numbers of syntactic frames, include various types of figuration, indicate an affective stance, and can describe situations of social interest (Nunberg et al., 1994).
To explore the capabilities and limitations of modern NLP systems for generating cross-lingual explanations of idiomatic expressions, we construct two datasets of English idioms and their explanations in English, Spanish, and German; EPIE-ME, an open dataset composed of 628 unique idioms with multi-lingual explanations built from the EPIE corpus (Saxena and Paul, 2020), and Oxford-ME, composed of 6,218 unique idioms taken from the Oxford Dictionary of Idioms 4th Edition (Ayto, 2020). We explore performance of both pre-trained sequence-to-sequence models and Large Language Models (LLMs), and show that our task configuration is much more challenging than previous definition generation tasks. We further provide some discussion on why naïve translation approaches do not suffice for this task.
@@
    We release a new cross-lingual dataset containing English idiomatic expressions as well as explanations of their meanings in English, Spanish, and German. [^1] Data and code available at: link upon publication. We are not able to publish the Oxford-ME portion due to copyright.
@@ 2 Related Work
Disambiguation of Non-Standard Language. Non-standard language takes many different forms ranging from slang to metaphor to idiom. To disambiguate some more general forms of non-standard language, Ni and Wang (2017) proposed a hybrid word-character sequence-to-sequence model that can disambiguate language found on social media. Kurfalı and Östling (2020) show that sufficiently different contextual embeddings can be found for both literal and idiomatic instances of a multi-word expression, thus allowing for disambiguation of the expressions in context. Qiang et al. (2023), on the other hand, try to directly paraphrase Chinese idioms such that they can replace the idiom using an infill strategy similar to the cloze task. These approaches are all tested in monolingual settings. The cross-lingual settings for which a solution might benefit language learners remain less explored.
@@
The cross-lingual version of idiomatic expression disambiguation has traditionally been approached as a translation task. Often, sentences containing idiomatic expressions are mapped to equivalent sentences without such an expression, but models may try to translate them word-by-word (Baziotis et al., 2023). Fadaee et al. (2018) explore sentence-level translation to generate an English–German idiom translation dataset, additionally using attention-based model to translate sentences containing idiomatic expressions, requiring in-context disambiguation. Zaninello and Birch (2020) show that data augmentation and external linguistic resources can sizably improve translations of multi-word expressions. Most recently, Li et al. (2024) take a very similar prompting-based approach to us. However, the authors’ primary goal appears to be the creation of an ontology, they perform their experiments with languages we do not consider, and they additionally remain focused on a translation-based perspective. We, however, formulate the task more closely to definition generation, and explore the capabilities of NLP systems for this purpose in addition to applications in high-stakes scenarios such as education. We believe that this reformulation opens up interesting research avenues, such as investigating whether end-users benefit more from explanations, examples usage, or a combination.
@@
Definition Generation. Since Noraset et al. (2017) introduced this task, new techniques have continued to be introduced to improve performance. Gadetsky et al. (2018) created embeddings from dictionary definitions, while taking into account the issue of polysemy by allowing their model to recognize that there could potentially be an infinite number of senses for a given word. In contrast, using a novel encoding scheme, Bevilacqua et al. (2020) are able to autoregressively generate definitions as well as score them in a discriminative manner. Zhang et al. (2023) generate cross-lingual definitions of words in English, Chinese, and French using prompt learning. They apply a contrastive learning objective to correct pre-trained models’ tendency to ignore this task and/or mix languages in the output. In this paper, we focus specifically on idiomatic expressions and their explanations, which allows us to be more flexible and we hypothesize is more challenging for NLP systems.
@@ 3.1 Task Definition
Our goal is to provide an explanation of the meaning of an idiomatic expression in their native language of a language learner with the assumption that it would be more understandable than an explanation in the language they are attempting to learn. Following prior work on definition generation, we formalize our task as a text-to-text generation task. Given an idiomatic expression $I=\{i_{1},i_{2},i_{3},...\}$ and some optional context $C=\{c_{1},c_{2},c_{3},...\}$ in the source language $L_{S}$; our goal is to produce an explanation $E=\{e_{1},e_{2},e_{3},...\}$ in the target language $L_{T}$. While this formulation is essentially identical to that of definition generation, we opt to call our generated text explanations. While minute, the difference we wish to highlight is that we feel a definition is generally more static as opposed to an explanation which may be tailored to an individual or particular context. Explanations might also include additional contextual information that can be realized in either speech or writing. While we do not require any of these factors to be present in our explanations, because our focus is on language learners and educational contexts, we wish to allow flexibility in how explanations are presented to a user in current and future exploration of this task.
@@
EN Avg Len 9.04 9.08 8.88
ES Avg Len 9.34 9.47 9.22
DE Avg Len 9.21 9.21 9.09
Table 1: Dataset Statistics
@@ 3.2 Dataset
We construct – to the best of our knowledge – the first two datasets for cross-lingual explanation generation of idiomatic expressions, EPIE-ME and Oxford-ME (EPIE and Oxford with Multilingual Explanations). We build on the EPIE corpus Saxena and Paul (2020), composed of 25,206 sentences containing instances of 717 formal and static idiomatic expressions with definitions, and example usages. [^2] Formal idiomatic expressions can undergo lexical changes, while static ones cannot. We select this dataset because it was in-part originally meant for detection of idiomatic expressions. As a result, it also contains expressions that are situationally idiomatic, a useful quality for judging explanation ability. We filter out expressions for which there are no gold explanations or in-context examples available, leaving us with 628. We additionally use the Oxford Dictionary of Idioms 4th Edition (Ayto, 2020), comprising of 6,218 English idiomatic expressions. The Oxford data only comes with sentence-level context for 1,780 of the expressions (28.6%), so we prompt Llama 3.1 to generate sentences using missing expressions in the manner found in Tab. 14. For each unique expression in the corpus, we include an explanation in English, Spanish, and German. We obtain non-English explanations by feeding the gold English explanations to Google Translate. In Tab. 1 we show the train/dev/test splits and the average word length of the idioms and explanations for both datasets. Splits are generated by random sample.
@@
Manual corrections. For the test sets, the English explanations and the automated translations are closely inspected and manually corrected by native speakers of all three languages to ensure they are entirely correct and not just the output of the automated translation. Any confusion about the meaning of a particular expression was referred to the native English speaker for review. These annotators are all computational linguists fluent in English and working in English speaking countries.
@@
Idiom categorization. For EPIE-ME, we further enhance our idioms with thematic categories inspired by findings suggesting that classifying idioms facilitates the learning process Boers (2000). To do this, we build on the taxonomy developed by Rafatbakhsh and Ahmadi (2019), which suggests a list of 81 themes to classify English idioms (See full list with frequencies in App. C). For any idioms already assigned themes by Rafatbakhsh and Ahmadi (2019) (a little over 20%), we use the assigned theme. To tag the remainder of our dataset using this taxonomy, we prompt GPT-3.5 Turbo to provide us with the three most appropriate categories over three runs where the categories are provided in random order. We then eliminate duplicates in the resulting nine predictions and correct them by hand. The manual corrections were done by two English speaking annotators working together to resolve conflicts.
@@
Estimating noise in non-corrected examples. Finally, we examine the amount of noise in the explanations without manual correction. We do so by calculating over our verified test set the edit distance (Levenshtein, 1965) between the original output of Google Translate and the corrected version produced by our annotators, and further normalize by length following Tashima et al. (2018). We see that the amount of correction required is generally low, the largest amount being required on average for the German EPIE-ME explanations with a normalized edit distance of 0.162. More information can be found in Appendix E.
@@ 4.1 Explanation Strategies
Direct Explanation. To obtain an explanation for the benefit of a potential language learner, we fine-tune and/or prompt models for an explanation in what we assume to be the L1 of the learner. All inputs, however, are in English. Our training pairs in the fine-tuning setup consist of an input $X_{i}=[T_{i,{L_{T}}};I_{i}]$ and label $Y_{i}=[E_{i,{L_{T}}}]$, where $T_{i,{L_{T}}}$ is set of tokens informing the model that its task is to produce an explanation in a target language $L_{T}$, $I_{i}$ is an idiomatic expression, and $E_{i,{L_{T}}}$ is a gold explanation of the idiomatic expression. This model can then be formulated as:
@@
Pipeline. In this configuration, we aim to generate our final explanation in two stages. The first stage is to generate an explanation in English. In our fine-tuning setup, we use T5 for this step. Where mT5 was explicitly designed to learn many languages, T5 was only trained to translate between a small set, and most of its training data was English. Thus, we feel T5 is more likely to perform well on this task. The model definition remains the same as in Eq. 1. However, the input pair is now $X_{i}=[T_{i};I_{i}]$ and $Y_{i}=[E_{i,L_{S}}]$. In this case, our task sequence $T_{i}$ no longer provides any language information, and our explanation $E_{i,L_{S}}$ remains in the source language (in our case, English).
@@
Table 2: Performance of Best Model on EPIE-ME Data per Theme for Top 10 Themes
@@ 4.2 Contextual Enhancements
Sentence-Level Context. Because idiomatic expressions encode information that is potentially unrelated to their linguistic form, many researchers have used sentence-level context to assist with identification, disambiguation, and translation of idiomatic expressions (Fadaee et al., 2018; Kurfali and Östling, 2020; Saxena and Paul, 2020; Liu et al., 2023; Li et al., 2024). We experiment with including example sentences (one usage example per instance) containing the idiomatic expression to be explained at the end of our prompts in an effort to provide more information to the model.
Categorical Enhancements. In this setting we explore how the addition of small hints can help these models. We add to our prompt a category that an idiomatic expression might belong to. All categories we use come from the Rafatbakhsh and Ahmadi (2019) taxonomy.
@@
Table 3: Best Model Results for Direct and Pipeline Strategies. LLM Results are Averaged Across 3 Runs on the Test Set. Scores for English Explanations in the First Half of the Pipeline can be Found in Tab. 8
@@ 4.3 Few-Shot Prompting
To induce a label, we ask the model to choose the most suitable label from our list of categories. If the induced label does not exactly match an existing label, we add it to our list and treat it as an existing label for only the instance in which we observe this idiomatic expression, and from this point forward, the procedure is the same as if we assume the category is known. Once we have a known category label (either given or acquired by induction), meaning we are given some idiomatic expression $I_{i}$ with category $C_{i}$, we form an initial selection of $2*k$ idiomatic expressions with category label $C_{j}$ where $C_{i}=C_{j}$ and $k$ is the number of shots. If there are too few expressions to form this selection, we draw from the most similar categories until we obtain $2*k$ examples. Similarity between categories is calculated as the similarity between the labels by the same SBERT model used for our evaluation in Sec. 5. Once these $2*k$ examples are obtained, we take a final sample of $k$ from this set. This sampling strategy is necessitated by the small size of our dataset, where we may not have even a single example in our training set with the same category label as one in our validation or test set. The first step of selecting $2*k$ initial examples helps reduce variance between instances where there are enough examples where the categories match and instances where there are too few.

### v3

@@ Abstract
Automated definition generation systems have been proposed to support vocabulary expansion for language learners. The main barrier to the success of these systems is that learners often struggle to understand definitions due to the presence of potentially unfamiliar words and grammar, particularly when non-standard language is involved. To address these challenges, we propose CLIX, the task of Cross-Lingual explanations of Idiomatic eXpressions. We explore the capabilities of current NLP models for this task and observe that, while it remains challenging, large language models show promise. Finally, we perform a detailed error analysis to highlight the key challenges that need to be addressed before we can reliably incorporate these systems into educational tools.
@@ 1 Introduction
Motivated by these challenges, we propose the task of Cross-Lingual explanation of Idiomatic eXpressions (CLIX). Given a source language idiomatic expression, we aim to generate a natural language explanation in a specified target language. As opposed to definition, we use the term explanation to allow for more flexibility in tailoring explanations to certain situations as well as potential inclusions of context. Rather than expecting a static, one-to-one mapping from idiom to definition, we allow models to approximate a one-to-many function. For example, explanations can be considered correct if they include elaborations, examples of use, and etymological information. The core concern of this research is the efficacy of the explanation and the final outcome for a potential student.
We additionally hypothesize that obtaining explanations of idiomatic expressions in the learner’s first language removes many of the barriers to understanding introduced by traditional definition generation systems. We choose to focus on idiomatic expressions as they are an important element of language learning that is particularly challenging for learners and automated systems alike. Consider the utterance, he and I don’t see eye to eye on a variety of topics. The idiomatic expression contained within this sentence is not composed of particularly challenging words, yet in this instance it denotes a negative stance relative to the referent’s opinions and its meaning has nothing to do with the general definitions of the words see and eye. This example helps to display the fact that idiomatic expressions often are of unpredictable form, appear in limited numbers of syntactic frames, include various types of figuration, indicate an affective stance, and can describe situations of social interest (Nunberg et al., 1994).
To explore the capabilities and limitations of modern NLP systems for generating cross-lingual explanations of idiomatic expressions, we construct two datasets of English idioms and their explanations in English, Spanish, and German: EPIE-ME, an open dataset composed of 628 unique idioms with multi-lingual explanations built from the EPIE corpus (Saxena and Paul, 2020), and Oxford-ME, composed of 6,218 unique idioms taken from the Oxford Dictionary of Idioms 4th Edition (Ayto, 2020). We explore performance of both pre-trained sequence-to-sequence models and Large Language Models (LLMs), and show that our task configuration is much more challenging than previous definition generation tasks. We further provide some discussion on why naïve translation approaches do not suffice for this task.
@@
    We release, to our knowledge, the first cross-lingual dataset for this task containing English idiomatic expressions as well as explanations of their meanings in English, Spanish, and German. [^1] Data and code available at: https://github.com/blast-cu/CLIX. We are not able to publish the Oxford-ME portion due to copyright, but can provide the data upon request.
@@ 2 Related Work
Non-standard language takes many different forms ranging from slang to metaphor to idiom. To disambiguate some more general forms of non-standard language, Ni and Wang (2017) proposed a hybrid word-character sequence-to-sequence model that can disambiguate language found on social media. Kurfalı and Östling (2020) show that sufficiently different contextual embeddings can be found for both literal and idiomatic instances of a multi-word expression, thus allowing for disambiguation of the expressions in context. Qiang et al. (2023), on the other hand, try to directly paraphrase Chinese idioms such that they can replace the idiom using an infill strategy similar to the cloze task. These approaches are all tested in monolingual settings. The cross-lingual settings for which a solution might benefit language learners remain less explored.
@@
The cross-lingual version of idiomatic expression disambiguation has traditionally been approached as a translation task. Often, sentences containing idiomatic expressions are mapped to equivalent sentences without such an expression, but models may try to translate them word-by-word (Baziotis et al., 2023). Fadaee et al. (2018) explore sentence-level translation to generate an English–German idiom translation dataset, additionally using attention-based model to translate sentences containing idiomatic expressions, requiring in-context disambiguation. Zaninello and Birch (2020) show that data augmentation and external linguistic resources can sizably improve translations of multi-word expressions. More recently, Li et al. (2024) take a very similar prompting-based approach to us. However, the authors’ primary goal appears to be the creation of an ontology, they perform their experiments with languages we do not consider, and they additionally remain focused on a translation-based perspective. We, however, formulate the task more closely to definition generation and explore the capabilities of NLP systems for this purpose in addition to applications in high-stakes scenarios such as education. We believe that this reformulation opens up interesting research avenues, such as investigating whether end-users benefit more from explanations, examples usage, or a combination. We also see new datasets specifically intended for determining whether an idiomatic expression is being used literally or figuratively given some context – now at the cross-lingual level (Sentsova et al., 2025). To the best of our knowledge, however, we are the first to release a cross-lingual dataset specifically for the purpose of explaining these kinds of expressions.
Definition Generation
@@
Since Noraset et al. (2017) introduced this task, new techniques have continued to be introduced to improve performance. Gadetsky et al. (2018) created embeddings from dictionary definitions, while taking into account the issue of polysemy by allowing their model to recognize that there could potentially be an infinite number of senses for a given word. In contrast, using a novel encoding scheme, Bevilacqua et al. (2020) are able to autoregressively generate definitions as well as score them in a discriminative manner. Zhang et al. (2023) generate cross-lingual definitions of words in English, Chinese, and French using prompt learning. They apply a contrastive learning objective to correct pre-trained models’ tendency to ignore this task and/or mix languages in the output. In this paper, we focus specifically on idiomatic expressions and their explanations, which allows us to be more flexible and we hypothesize is more challenging for NLP systems.
@@ 3.1 Task Definition
Our goal is to provide an explanation of the meaning of an idiomatic expression in their native language of a language learner with the assumption that it would be more understandable than an explanation in the language they are attempting to learn. Following prior work on definition generation, we formalize our task as a text-to-text generation task. Given an idiomatic expression $I=\{i_{1},i_{2},i_{3},...\}$ and some optional context $C=\{c_{1},c_{2},c_{3},...\}$ in the source language $L_{S}$, our goal is to produce an explanation $E=\{e_{1},e_{2},e_{3},...\}$ in the target language $L_{T}$. While this formulation is essentially identical to that of definition generation, we opt to call our generated text explanations. While minute, the difference we wish to highlight is that we feel a definition is generally more static as opposed to an explanation which may be tailored to an individual or particular context. Explanations might also include additional contextual information that can be realized in either speech or writing. While we do not require any of these factors to be present in our explanations, because our focus is on language learners and educational contexts, we wish to allow flexibility in how explanations are presented to a user in current and future exploration of this task. Furthermore, EPIE (Saxena and Paul, 2020) and our modified version EPIE-ME do not only contain static definitions. We see samples such as cheek and jowl which are instead given an explanation: Positioned very close together. The cheek and the jowl — the lower part of the jaw — are in close proximity to each other on the face.
@@
EN Avg Len 9.04 9.08 9.08
ES Avg Len 9.34 9.47 9.11
DE Avg Len 9.21 9.21 9.06
Table 1: Dataset Statistics. I Refers to Idiom, EN, ES, and DE Refer to English, Spanish, and German Explanations
@@ 3.2 Dataset
We construct – to the best of our knowledge – the first two datasets for cross-lingual explanation generation of idiomatic expressions, EPIE-ME and Oxford-ME (EPIE and Oxford with Multilingual Explanations), each containing explanations from languages belonging to two separate branches of the Indo-European family, Spanish and German. We build on the EPIE corpus Saxena and Paul (2020), composed of 25,206 sentences containing instances of 717 formal and static idiomatic expressions with definitions, and example usages. [^2] Formal idiomatic expressions can undergo lexical changes, while static ones cannot. We select this dataset because it was in-part originally meant for detection of idiomatic expressions. As a result, it also contains expressions that are situationally idiomatic, a useful quality for judging explanation ability. We filter out expressions for which there are no gold explanations or in-context examples available, leaving us with 628. We additionally use the Oxford Dictionary of Idioms, 4th Edition (Ayto, 2020), comprising of 6,218 English idiomatic expressions and their explanations. This is generally considered the most comprehensive resource for idiomatic expressions in English. The Oxford data only comes with sentence-level context for 1,780 of the expressions (28.6%), so we prompt Llama 3.1 (Dubey et al., 2024) to generate sentences using missing expressions in the manner found in Tab. 14. For each unique expression in the corpus, we include an explanation in English, Spanish, and German. We obtain non-English explanations by feeding the gold English explanations to Google Translate. In Tab. 1 we show the train/dev/test splits and the average word length of the idioms and explanations for both datasets. Splits are generated by random sample.
Manual Corrections
@@
For the test sets, the English explanations and the automated translations are closely inspected and manually corrected by native speakers of all three languages to ensure they are entirely correct and not just the output of the automated translation. Any confusion about the meaning of a particular expression was referred to the native English speaker for review. These annotators are all computational linguists fluent in English and working in English speaking countries.
Idiom Categorization
@@
For EPIE-ME, we further enhance our idioms with thematic categories inspired by findings suggesting that classifying idioms facilitates the learning process Boers (2000). To do this, we build on the taxonomy developed by Rafatbakhsh and Ahmadi (2019), which suggests a list of 81 themes to classify English idioms (See full list with frequencies in App. C). For any idioms already assigned themes by Rafatbakhsh and Ahmadi (2019) (a little over 20%), we use the assigned theme. To tag the remainder of our dataset using this taxonomy, we prompt GPT-3.5 Turbo (OpenAI, 2023) to provide us with the three most appropriate categories over three runs where the categories are provided in random order. We then eliminate duplicates in the resulting nine predictions and correct them by hand. The manual corrections were done by two English speaking annotators working together to resolve conflicts.
Estimating Noise in Non-Corrected Examples
@@
Finally, we examine the amount of noise in the explanations without manual correction. We do so by calculating the edit distance (Levenshtein, 1965) between the original output of Google Translate and the corrected version produced by our annotators over our verified test set. We further normalize by length following Tashima et al. (2018). We see that the amount of correction required is generally low, the largest amount being required on average for the German EPIE-ME explanations with a normalized edit distance of 0.283, while Spanish had an edit distance of 0.159. More information can be found in Appendix E.
@@ 4.1 Explanation Strategies
Direct Explanation
To obtain an explanation for the benefit of a potential language learner, we fine-tune and/or prompt models for an explanation in what we assume to be the L1 of the learner. All inputs, however, are in English. Our training pairs in the fine-tuning setup consist of an input $X_{i}=[T_{i,{L_{T}}};I_{i}]$ and label $Y_{i}=[E_{i,{L_{T}}}]$, where $T_{i,{L_{T}}}$ is set of tokens informing the model that its task is to produce an explanation in a target language $L_{T}$, $I_{i}$ is an idiomatic expression, and $E_{i,{L_{T}}}$ is a gold explanation of the idiomatic expression. This model can then be formulated as:
@@
Pipeline
In this configuration, we aim to generate our final explanation in two stages. The first stage is to generate an explanation in English. In our fine-tuning setup, we use T5 (Raffel et al., 2020) for this step. Where mT5 (Xue et al., 2021) was explicitly designed to learn many languages, T5 was only trained to translate between a small set, and most of its training data was English. Thus, we feel T5 is more likely to perform well on this task. The model definition remains the same as in Eq. 1. However, the input pair is now $X_{i}=[T_{i};I_{i}]$ and $Y_{i}=[E_{i,L_{S}}]$. In this case, our task sequence $T_{i}$ no longer provides any language information, and our explanation $E_{i,L_{S}}$ remains in the source language (in our case, English).
@@
Table 2: Performance of Best Model as Measured by Sentence Similarity on EPIE-ME Data per Theme for Top 10 Themes
@@ 4.2 Contextual Enhancements
Sentence-Level Context
Because idiomatic expressions encode information that is potentially unrelated to their linguistic form, many researchers have used sentence-level context to assist with identification, disambiguation, and translation of idiomatic expressions (Fadaee et al., 2018; Kurfali and Östling, 2020; Saxena and Paul, 2020; Liu et al., 2023; Li et al., 2024). We experiment with including example sentences (one usage example per instance) containing the idiomatic expression to be explained at the end of our prompts in an effort to provide more information to the model.
Categorical Enhancements
In this setting we explore how the addition of small hints can help these models. We add to our prompt a category that an idiomatic expression might belong to. All categories we use come from the Rafatbakhsh and Ahmadi (2019) taxonomy.
@@
Table 3: Best Model Results by Sentence Similarity for Direct and Pipeline Strategies. LLM Results are Averaged Across 3 Runs on the Test Set. Scores for English Explanations in the First Half of the Pipeline can be Found in Tab. 8
@@ 4.3 Few-Shot Prompting
To induce a label, we ask the model to choose the most suitable label from our list of categories. If the induced label does not exactly match an existing label, we add it to our list and treat it as an existing label for only the instance in which we observe this idiomatic expression, and from this point forward, the procedure is the same as if we assume the category is known. Once we have a known category label (either given or acquired by induction), meaning we are given some idiomatic expression $I_{i}$ with category $C_{i}$, we form an initial selection of $2*k$ idiomatic expressions with category label $C_{j}$ where $C_{i}=C_{j}$ and $k$ is the number of shots. If there are too few expressions to form this selection, we draw from the most similar categories until we obtain $2*k$ examples. Similarity between categories is calculated as the similarity between the labels by the same SBERT (Reimers and Gurevych, 2020) model used for our evaluation in Sec. 5. Once these $2*k$ examples are obtained, we take a final sample of $k$ from this set. This sampling strategy is necessitated by the small size of our dataset, where we may not have even a single example in our training set with the same category label as one in our validation or test set. The first step of selecting $2*k$ initial examples helps reduce variance between instances where there are enough examples where the categories match and instances where there are too few.